OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x A. 2 sin x - sin3x - cos x B. 2 sin x cos2x + sin x - 2 sin3x - cos x C.2 sin3x cos4x + 1 D. 3 sin x cos x - sin3x - cos x

OpenStudy (anonymous):

PLEASE HELP ANYBODY :)

OpenStudy (mertsj):

\[\sin(3x)-cosx=\sin(2x+x)-cosx=\sin2xcosx+\cos2xsinx-cosx\]

OpenStudy (mertsj):

\[2sinxcosxcosx+(\cos^2x-\sin^2x)sinx -cosx=2sinxcos^2x+\cos^2x-\sin^3x-cosx\]

OpenStudy (mertsj):

Is the original problem sin(3x) or sin^3x?

OpenStudy (anonymous):

i am not sure,, what i typed is what was given

OpenStudy (anonymous):

sin 3x-cos x

OpenStudy (mertsj):

It's D

OpenStudy (anonymous):

thank you so much! !!!

OpenStudy (mertsj):

yw